Perform the following steps to verify by substitut
Answer #1


dE/dx = -kEmax sin(kx -wt)

d^2E/dx^2 = -k^2 Emax cos(kx -wt)

dB/dx = -kBmax sin(kx -wt)

d^2B/dx^2 = -k^2 Bmax cos(kx -wt)

dE/dt = wEmax sin(kx -wt)

d^2E/dt^2 = -w^2 Emax cos(kx -wt)

dB/dt = wBmax sin(kx -wt)

d^2B/dt^2 = -w^2 Bmax cos(kx -wt)


(d^2E/dx^2)/(d^2E/dt^2) = k^2/w^2 = (1/f*lamda)^2 = 1/c^2

(d^2E/dx^2)/(d^2E/dt^2) = (1/c)^2

(d^2B/dx^2)/(d^2B/dt^2) = k^2/w^2

(d^2B/dx^2)/(d^2B/dt^2) = (1/c)^2

(d) eo*uo = = (1/c)^2

Here (1/c)^2 = 1.11*10^-17 s^2/m^2